OpenStudy (anonymous):

how do I prove (1-sin^4(theta))/(1+sin^2(theta)) =cos^2(theta)

OpenStudy (freckles):

factor the numerator on the left hand side (think difference of squares)

OpenStudy (anonymous):

(1-sin^2(theta))(1+sin^2(theta))

OpenStudy (freckles):

right! and you see something cancels

OpenStudy (anonymous):

then I cancel out and use which Identity

OpenStudy (freckles):

\[\frac{(1-\sin^2(x))\cancel{(1+\sin^2(x))}}{\cancel{1+\sin^2(x)}}\] think Pythagorean identity

OpenStudy (anonymous):

yes you are right I just could not figure it out but I forgot that any of the steps in the Pythagorean make it right Thanks so much!

OpenStudy (freckles):

right the one you want is the one that says sin^2(x)+cos^2(x)=1

OpenStudy (anonymous):

yes Thanks Freckles

OpenStudy (freckles):

np